OpenStudy (logan13):

A square painting has an area of 1040 in2. To the nearest tenth, what is the length of 1 side of the painting?

OpenStudy (logan13):

@texaschic101

OpenStudy (mathstudent55):

The area of a square with side x is \(A = x^2\) You have the area, so you have \(x^2 = 1040 ~in^2\) All you need to do is take the square root of 1040 to find x, the side. Use a calculator to find the square root of 1040. Then round off to the nearest tenth.

OpenStudy (mahoganie.carson):

The area of square = (sidelength)^2 1040 = L^2 L = rootof(1040) L= 32.2 Here you go hope this helps :)

OpenStudy (mahoganie.carson):

@logan13